energy released in fission of uranium is 200Mev where as enrgy released in fusion of 4 hydrogen nuclei is 24.7Mev which is greater

Grade:12th Pass

1 Answers

Khimraj
6 years ago
energy released in fusion is greater because  energy released always compared in per unit mass.
energy reased in fission of uranium = 200/238 = 0.84Mev/amu
energy reased in fusionof 4 hydrogen = 24/4 = 6Mev/amu.
Hope it clears. If still you doubt you can ask me.
If you like anser then please approve it.